OpenStudy (anonymous):

If a line with a slope of -2 crosses the y-axis at (0, 3), what is the equation of the line? y − 2x = 3 -2y + 3x = 1 y + 3x = 2 y + 2x = 3

OpenStudy (anonymous):

We can write the equation of a line as: (y-b) = m (x-a) where m = the slope of the line, and (a,b) is a point on the line. Substitute your values of m, a, and b into the above equation

OpenStudy (anonymous):

Refer to the attachment.